OpenStudy (anonymous):

Which statement has the solution shown on the number line? |x – 2| = 1 |2x – 2| = 6 |3x – 3| = 6 none of the above ON THE NUMBER LINE, THE NUMBERS -1 and 3 are both marked.

OpenStudy (jamesj):

Let's take these each in turn. |x-2| = 1 gives all the numbers that are a distance of 1 unit from 2. Agreed? What are those two numbers?

OpenStudy (anonymous):

wait what am i supposed to do?

OpenStudy (jamesj):

I'm helping you understand what are the solutions for each of these equations so you can compare them with the given solutions of -1 and 3. If we find an equation which has solutions -1 and 3, then that equation is the answer to your problem.

OpenStudy (jamesj):

So the first question is: what are the solutions to |x-2| = 1 ?

OpenStudy (anonymous):

do i plug in -1 and -3 in there or what?

OpenStudy (jamesj):

No, let's not go down that road. Geometrically |x-2| = 1 means all numbers x which are at a distance of 1 away from 2. What numbers are 1 away from 2? 1 and 3. 1 and 3 are the solutions to this equation. Hence this is not the equation you're looking for.

OpenStudy (jamesj):

Second equation |2x – 2| = 6 Divide both sides by 2 and you get |x-1| = 3 What are the solutions of that equation? You want numbers that are a distance 3 away from 1.

OpenStudy (anonymous):

so the solutions are -1 and 3?

OpenStudy (jamesj):

No, 3 away from 1 is either -2 or 5. Do you see that?

OpenStudy (anonymous):

i dont see how its 5? shouldn't it be 4?

OpenStudy (jamesj):

need coffee! Yes, sorry, -2 or 4.

OpenStudy (jamesj):

Third equation |3x – 3| = 6 Divide both sides by 3 |x-1| = 2 What are the solutions of this equation?

OpenStudy (anonymous):

ITS NOT -2 EITHER ITS -4

OpenStudy (jamesj):

The solutions of |x-1| = 3 are 1-3 = -2 and 1+3 = 4

OpenStudy (anonymous):

ok so which one of the choices do you think it will be?

OpenStudy (jamesj):

Third equation |3x – 3| = 6 Divide both sides by 3 |x-1| = 2 What are the solutions of this equation?

OpenStudy (anonymous):

HOW DO I FIND A SOLUTION?!?!?!?!

OpenStudy (jamesj):

|x-1|=2 means numbers that are a distance 2 away from 1. hence the solutions are 1-2 and 1+2 which are ... what?

OpenStudy (anonymous):

so which one of the choices is it?

OpenStudy (jamesj):

You're not following. What are the solutions to the third equation |3x – 3| = 6

OpenStudy (jamesj):

The solutions to that equation are the solutions to |x-1| = 2 and the solutions to that equation are what?

OpenStudy (anonymous):

I DONT KNOW!!! Ok, listen ive been working on this ALL day yesterday till 2am, got waken up today at 5am just to finish this. ive had enough. i have like no more brain power, and i dont have time for this. im bout to reach the deadline. so uhh, thanks for the help? i guess.

OpenStudy (jamesj):

You are one step away from solving the problem. |x-1| = 2 is the set of numbers which are a distance 2 away from 1. Those two numbers are 1+2 = 3 and 1-2 = -1 ---------- See the answer to your entire problem now?

OpenStudy (anonymous):

no

OpenStudy (jamesj):

You are told the solution to a particular equation is two numbers: -1 and 3. The task is to find the equation which has -1 and 3 as its solution. Didn't we just find the equation which has -1 and 3 as its solution?

OpenStudy (jamesj):

Still there?